# Searching Accretion-Enhanced Dark Matter Annihilation Signals in the Galactic Centre

### v1

@@ I Introduction
In this study, we revisit the accretion-enhanced annihilation process in a DM halo spike, with a focus on a fermionic DM model incorporating an electroweak axion-like particle (ALP) portal Peccei and Quinn (1977a, b); Weinberg (1978); Wilczek (1978a). We explore three specific annihilation channels (new mediator resonance, $p$-wave secluded DM and forbidden DM) based on mass conditions, analyzing the photon yield spectrum and annihilation cross-section relative to the position and velocity of DM particles with respect to the SMBH. For gamma-ray continuum spectra, we exclusively utilize Fermi data to establish upper limits on the ALP-DM-DM coupling $C_{a\chi\chi}$ across three scenarios. Additionally, following the methodology developed in Ref. Cheng et al. (2023a), we combine publicly available 6 years of DAMPE and 15 years of Fermi-LAT data from Sgr A to search for spectral line signals.
@@ II The DM model with electroweak ALP portal
The annihilation rates of DM to photons are determined by the couplings $C_{a\gamma\gamma}$ and $C_{a\chi\chi}$.
Generally speaking, DM annihilations to photons in this model can be dominant by the $s$-wave contribution via $a$-resonance, or $p$-wave contribution ($\chi\overline{\chi}\to aa$). In addition, some different mass combinations between $\chi$ and $a$ can alter the velocity dependence of the DM annihilation cross-section. If DM particles can be accelerated and annihilated to photons as presented in Fig. 1, three possible annihilation mechanisms based on the relationships between $m_{\chi}$ and $m_{a}$ appear as what follows,
@@
For simplicity, we use the following assumptions for the model parameters in our numerical studies without loss of generality. First, we set $C_{WW}$ = $C_{BB}$ to reduce one free model parameter. Then, the electroweak ALP strongly suffers from collider constraints Allen et al. (2024): $C_{a\gamma\gamma}\lesssim 2.5\times 10^{-5}\;\rm GeV^{-1}$ for $m_{a}\sim 1$ TeV from the LHC ALP to diphoton searches Jaeckel et al. (2013), and $C_{a\gamma\gamma}\lesssim 1.0\times 10^{-6}\;\rm GeV^{-1}$ for $m_{a}\sim 4$ GeV from the BaBar $B\to K\gamma\gamma$ searches Lees et al. (2022). We will see latter that cross-sections for $p$-wave, forbidden annihilation and $a$-resonance are not sensitive to the size of ALP couplings to the SM gauge bosons. Therefore, we can simply set $C_{a\gamma\gamma}=2.5\times 10^{-5}\;\rm GeV^{-1}$ in our numerical calculations.
@@ III DM annihilation in the spike halo
where $\rho_{s}\approx 0.0156\;\rm M_{\odot}/\rm pc^{3}$ and $r_{s}=1.5\times 10^{4}\;\rm pc$ are the scaled density and scaled radius McMillan (2016). Here, we take the slope of the density profile $\gamma=0.8$ based on the S-star combined constraints $\gamma<0.83$ Shen et al. (2023). The DM-induced gravitational potential with positive-definition can be written as
[…]
@@
We can utilize the parametrization of $I_{r}^{\prime}$ provided in Gondolo and Silk (1999), which relies on the power-law profile and exhibits uncertainties below $8\%$, to establish a connection with our expression in Eq. (8), allowing us to derive the relationship among $\epsilon$, $\epsilon^{\prime}$, and $L^{\prime}$. Once we obtain $\epsilon(\epsilon^{\prime},L^{\prime})$ and $f^{\prime}(\epsilon^{\prime},L^{\prime})=f_{0}\left[\epsilon(\epsilon^{\prime},L^{\prime})\right]$, the spike halo profile can be calculated by integrating the final-state phase space,
@@
Figure 2: The halo density profile (left panel) and gravitational potential (right panel) as a function of radius. The gray region $r<2R_{s}$ represents a no-DM particle zone. The spike halo profile (red solid line) represents the collisionless DM paradigm. The blue lines represent velocity-averaged annihilation cross-sections with $m_{\chi}=1\,\mathrm{GeV}$: $\langle\sigma v_{\rm rel.}\rangle=10^{-26}$ cm3s-1 (solid) and $\langle\sigma v_{\rm rel.}\rangle=10^{-30}$ cm3s-1 (dashed). The spike halo profile matches the NFW profile (depicted as a black solid line) at $r\geq r_{\rm sp}$. As a comparison in the right panel, the total gravitational potential $\Phi_{\rm tot}$ (illustrated as a red dashed line) contains contributions from both BH (black dashed line) and DM (green dashed line).
@@
In the right panel of Fig. 2, the potential functions from BH, DM, and the total potential function are represented by black, green, and red dot-dashed lines. Within the innermost region, BH gravitational potential dominates, accelerating DM to half the speed of light at $2R_{s}$. Once the radius exceeds $r_{\rm sp}$, the DM spike potential becomes dominant, naturally causing the DM profile to recover to the NFW profile.
@@
${\rm with}\quad\mu_{0}\equiv$ $\frac{4v_{\rm esc}^{2}-4v_{\rm CM}^{2}-v_{\rm rel.}^{2}}{4v_{\rm rel.}v_{\rm CM}},\quad\cos\alpha\equiv\frac{\mathbf{v_{\rm CM}}\cdot\mathbf{r}}{v_{\rm CM}\times r}\quad{\rm and}\quad v_{\rm esc}=\sqrt{2\Phi_{\rm tot}(r)}.$
@@
With a velocity condition $v>v_{\rm esc}[r_{\rm ann}]$ indicating their locations within $r<r_{\rm ann}$, such DM particles at $r=10^{-6}\,\mathrm{pc}$ (red dash-dotted line) can fully annihilate. In contrast, the high-speed tail of collisionless DM particles can still be truncated at $v_{\rm esc}$, but there is no annihilation to consume high-speed DM particles. Thus, weak DM annihilation facilitates the accumulation of high-speed DM particles. However, there is almost no difference made by DM annihilation in $r>r_{\rm ann}$, as seen from the overlapping blue dash-dotted and blue solid lines ($r=10^{-2}\,\mathrm{pc}$) in Fig. 3.
@@ III.1 Scenario $m_{\chi}>m_{a}$
In Fig. 4, the left panel shows results of $\langle\sigma v_{\rm rel.}\rangle$ with $C_{a\chi\chi}=3\times 10^{-5}/\,\mathrm{GeV}$ for $p$-wave annihilation, while the right panel presents $\langle dN_{\gamma}/dE_{\gamma}\rangle[r]$. Based on the right panel of Fig. 2, we observe that $\Phi\propto r^{-1}$ or $v_{\rm esc}\propto r^{-1/2}$ for $r<10^{2}\,\mathrm{pc}$ where $v_{\rm esc}$ increases with the radius. Together with $\sigma v_{\rm rel.}\propto v_{\rm rel.}^{2}$ for $p$-wave feature, $\langle\sigma v_{\rm rel.}\rangle$ in the left panel of Fig. 4 follows $\sigma v_{\rm rel.}\propto r^{-1}$ for $r<10^{2}\,\mathrm{pc}$, while it increases with radius for $r>10^{2}\,\mathrm{pc}$. In addition, $\langle\sigma v_{\rm rel.}\rangle$ increases with $m_{\chi}^{2}$ for $m_{\chi}\gg m_{a}$ as presented in Eq. (18), leading to dash lines ($m_{\chi}=250\,\mathrm{GeV}$) larger than solid lines ($m_{\chi}=25\,\mathrm{GeV}$) by two orders of magnitude. Similarly, those red lines ($m_{a}/m_{\chi}=0.9$) are one order of magnitude smaller than the blue lines ($m_{a}/m_{\chi}=0.1$), because a larger value of $m_{a}/m_{\chi}$ corresponds to a smaller $\langle\sigma v_{\rm rel.}\rangle$.
@@ III.2 Resonance annihilation
where $M_{Z}$ and $M_{W}$ are the masses of the $Z$ and $W$ bosons, respectively. With the assumption of $C_{BB}=C_{WW}$, the partial decay widths of $a\to\chi\overline{\chi}$, $a\to\gamma\gamma$, $a\to ZZ$, and $a\to W^{+}W^{-}$ can be found in detail in Appendix C. The decay mode $a\to\gamma Z$ vanished when $C_{BB}=C_{WW}$.
@@ IV Gamma-Ray Flux from GC
Similarly, in the forbidden DM scenario, BP4 (gold), BP5 (pink), and BP6 (cyan) exhibit spectra in the lower panels. For the forbidden DM scenario with $m_{\chi}=50\,\mathrm{GeV}$, BP4 and BP5 show continuum spectra, whereas BP6 predicts a spectral line, with parameters:
@@ V.2 Constraints of sharp spectral structures from the combined Fermi-LAT and DAMPE data
Thanks to the large statistics, the Fermi-LAT data are widely used in the search of the line-like (Abdo et al., 2010; Ackermann et al., 2015; Bringmann et al., 2012; Liang et al., 2016; Foster et al., 2023; De La Torre Luque et al., 2024) and box-like (Ibarra et al., 2012; Li et al., 2018; Shen et al., 2021) structures. We use the 15-yr P8R3_V3 CLEAN Fermi-LAT data (Ajello et al., 2021) between 2008 August 4 and 2023 August 3 (Fermi Mission Elapsed Time (MET) from 239557417 to 712715727) to search for sharp spectral structures. To improve the sensitivity of sharp structures, we only adopt the evtype=896 events, which removes $\approx 1/4$ of the data with the worst reconstructed energy quality. The same zenith cut and data-quality filter cut as the previous subsection are applied with Fermitools.
@@ VII Summary and Discussion
Finally, we aim to compare our findings with existing researches in the field. Previous studies Chiang et al. (2020); Alvarez and Yu (2021); Liu et al. (2022); Balaji et al. (2023) focus on $s$-wave annihilation. In contrast, only $a$-resonance annihilation in our work is under $s$-wave condition, highlighting scenario where annihilation cross-sections are enhanced near the BH by reaching the pole velocity. Ref. Cheng et al. (2023b) also examines resonance annihilation through a Higgs portal DM model, while annihilation into photon pairs is not the tree-level contribution. Some studies investigate $p$-wave annihilation near the Sgr A* Shelton et al. (2015); Sandick et al. (2018); Johnson et al. (2019); Cheng et al. (2022) and in the center of M87 Christy et al. (2023). Only Ref. Johnson et al. (2019) examines the axion-portal DM model, but we derive constraints using the halo slope given by S2 data rather than $\gamma\geq 1$. In Appendix D, we show the impact of $\gamma$ on spike halo density and the upper limits of $C_{a\chi\chi}$ in the $a$-resonance scenario. Furthermore, Ref. Johnson et al. (2019) claims that constraints derived from the continuum analyses are more stringent than those from line analyses. However, our analysis demonstrates that integrating both line and continuum analyses using data from Fermi and DAMPE can offer complementary limits. Ref. Shelton et al. (2015) also discusses various search strategies for lines or continuum signals depending on parameter space, but they focus on the Higgs-portal model. In contrast, we demonstrate that the axion-portal model supports a monochromatic line in both $a$-resonance and forbidden DM scenarios.
@@ Acknowledgments
YST, MWY, ZQG and ZQX are supported by the National Key Research and Development Program of China (No. 2022YFF0503304), and the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). CTL and XYL are supported by the Special funds for postdoctoral overseas recruitment, Ministry of Education of China (No. 164080H0262403). This research has made use of data obtained from the High Energy Astrophysics Science Archive Research Center (HEASARC), provided by NASA’s Goddard Space Flight Center. This research has made use of the data resources from DArk Matter Particle Explorer (DAMPE) satellite mission supported by Strategic Priority Program on Space Science, and China data service provided by National Space Science Data Center of China.
Appendix A DM self-interaction

### v2

@@ I Introduction
In this study, we revisit the accretion-enhanced annihilation process in a DM halo spike, with a focus on a fermionic DM model incorporating an electroweak axion-like particle (ALP) portal Peccei and Quinn (1977a, b); Weinberg (1978); Wilczek (1978a); Ghosh et al. (2024). We explore three specific annihilation channels (new mediator resonance, $p$-wave secluded DM and forbidden DM) based on mass conditions, analyzing the photon yield spectrum and annihilation cross-section relative to the position and velocity of DM particles with respect to the SMBH. For gamma-ray continuum spectra, we exclusively utilize Fermi data to establish upper limits on the ALP-DM-DM coupling $C_{a\chi\chi}$ across three scenarios. Additionally, following the methodology developed in Ref. Cheng et al. (2023a), we combine publicly available 6 years of DAMPE and 15 years of Fermi-LAT data from Sgr A to search for spectral line signals.
@@ II The DM model with electroweak ALP portal
The annihilation rates of DM to photons are determined by the couplings $C_{a\gamma\gamma}$ and $C_{a\chi\chi}$. For simplicity, we use the following assumptions for the model parameters in our numerical studies without loss of generality. First, we set $C_{WW}$ = $C_{BB}$ to reduce one free model parameter. Then, the electroweak ALP strongly suffers from collider constraints Allen et al. (2024): $C_{a\gamma\gamma}\lesssim 2.5\times 10^{-5}\;\rm GeV^{-1}$ for $m_{a}\sim 1$ TeV from the LHC ALP to diphoton searches Jaeckel et al. (2013), and $C_{a\gamma\gamma}\lesssim 1.0\times 10^{-6}\;\rm GeV^{-1}$ for $m_{a}\sim 4$ GeV from the BaBar $B\to K\gamma\gamma$ searches Lees et al. (2022).
Generally speaking, DM annihilations to photons in this model can be dominant by the $s$-wave contribution via $a$-resonance, or $p$-wave contribution ($\chi\overline{\chi}\to aa$). In addition, some different mass combinations between $\chi$ and $a$ can alter the velocity dependence of the DM annihilation cross-section. To highlight various velocity-dependent features of the annihilation cross section, we set $C_{a\gamma\gamma}=10^{-9}\,\mathrm{GeV}^{-1}$ for $m_{a}>500\;\rm MeV$ in numerical calculations when $m_{\chi}>m_{a}$ and $m_{\chi}\lesssim m_{a}$, thereby neglecting the $s$-channel contribution. In these cases, the main processes involve $p$-wave annihilation (velocity-squared suppression) and forbidden annihilation (velocity-exponential suppression). When $2m_{\chi}\approx m_{a}$, the decay channel $a\to\chi\chi$ opens, we assign $C_{a\gamma\gamma}=2.5\times 10^{-5}\,\mathrm{GeV}^{-1}$ to increase the efficiency of DM annihilation to photons.
If DM particles can be accelerated and annihilated to photons as presented in Fig. 1, three possible annihilation mechanisms based on the relationships between $m_{\chi}$ and $m_{a}$ appear as what follows,
@@ III DM annihilation in the spike halo
where $\rho_{s}\approx 0.0156\;\rm M_{\odot}/\rm pc^{3}$ and $r_{s}=1.5\times 10^{4}\;\rm pc$ are the scaled density and scaled radius McMillan (2016). Here, we take the slope of the density profile $\gamma=0.8$ based on the S-star combined constraints $\gamma<0.83$ Shen et al. (2023). The DM-induced gravitational potential with positive definition can be written as Yang and Yu (2022)
$\Phi(r)=4\pi{\rm G}\left[\frac{1}{r}\int_{0}^{r}\rho_{\rm NFW}(r^{\prime})r^{\prime 2}{\rm d}r^{\prime}+\int_{r}^{\infty}\rho_{\rm NFW}(r^{\prime})r^{\prime}{\rm d}r^{\prime}\right].$ (5)
@@
We can utilize the parametrization of $I_{r}$ provided in Gondolo and Silk (1999), which relies on the power-law profile and exhibits uncertainties below $8\%$, to establish a connection with our expression in Eq. (8), allowing us to derive the relationship among $\epsilon$, $\epsilon^{\prime}$, and $L^{\prime}$. Once we obtain $\epsilon(\epsilon^{\prime},L^{\prime})$ and $f^{\prime}(\epsilon^{\prime},L^{\prime})=f_{0}\left[\epsilon(\epsilon^{\prime},L^{\prime})\right]$, the spike halo profile can be calculated by integrating the final-state phase space,
@@
Figure 2: The halo density profile (left panel) and the rotation velocity squared (right panel) as a function of radius. The gray region $r<2R_{s}$ represents a no-DM particle zone. The spike halo profile (red solid line) represents the collisionless DM paradigm. The blue lines represent velocity-averaged annihilation cross-sections with $m_{\chi}=1\,\mathrm{GeV}$ for $\langle\sigma v_{\rm rel.}\rangle=10^{-26}$ cm3s-1 (solid) and $\langle\sigma v_{\rm rel.}\rangle=10^{-30}$ cm3s-1 (dot-dashed). The spike halo profile matches the NFW profile (depicted as a black solid line) at $r\geq r_{\rm sp}$. As a comparison in the right panel, the total rotation velocity squared (illustrated as a red dot-dashed line) contains the contributions from both BH (black dot-dashed line) and DM (green dot-dashed line).
@@
In the right panel of Fig. 2, the rotation velocity squared
[…]
The individual contributions to $v_{\rm rot}^{2}$ due to gravity are BH (black line), DM (green line), and the sum (red dot-dashed line). Within the innermost region, BH gravity dominates and accelerates DM particles to half the speed of light at $2R_{s}$. When $r>r_{\rm sp}$, the gravity of the DM halo dominates, leading the density profile to revert to the NFW.
@@
${\rm with}\quad\mu_{0}\equiv$ $\frac{4v_{\rm esc}^{2}-4v_{\rm CM}^{2}-v_{\rm rel.}^{2}}{4v_{\rm rel.}v_{\rm CM}},\quad\cos\alpha\equiv\frac{\mathbf{v_{\rm CM}}\cdot\mathbf{r}}{v_{\rm CM}\times r}\quad{\rm and}\quad v_{\rm esc}=\sqrt{2\Phi_{\rm total}(r)},$
where the graph of the potential function $\Phi_{\rm total}(r)$ is displayed in Appendix A.
@@
For the benchmark (dash-dotted lines) shown in Fig. 3, annihilation mainly occurs within $r\leq r_{\rm ann}\approx 10^{-3}\text{ pc}$, where the velocities exceed $v_{\rm esc}[r_{\rm ann}]$ (the escape velocity at $r_{\rm ann}$) and significantly enhance the velocity-dependent annihilation rate. Thus, at a smaller radius $r=10^{-6}\text{ pc}$, DM particles with $v_{\rm esc}[r_{\rm ann}]\leq v<v_{\rm esc}[10^{-6}\text{ pc}]$ can fully annihilate, leading to a truncation of the high-velocity tail of the velocity distribution, as illustrated by the leftward shift of the peak position in Fig. 3. In contrast, the high-speed tail of collisionless DM particles can still be truncated at $v_{\rm esc}$, but there is no annihilation to consume high-speed DM particles. Thus, weak DM annihilation facilitates the accumulation of high-speed DM particles. However, there is almost no difference made by DM annihilation in $r>r_{\rm ann}$, as seen from the overlapping blue dash-dotted and blue solid lines ($r=10^{-2}\,\mathrm{pc}$) in Fig. 3.
@@ III.1 Scenario $m_{\chi}>m_{a}$
In Fig. 4, the left panel shows results of $\langle\sigma v_{\rm rel.}\rangle$ with $C_{a\chi\chi}=3\times 10^{-5}/\,\mathrm{GeV}$ for $p$-wave annihilation, while the right panel presents $\langle dN_{\gamma}/dE_{\gamma}\rangle[r]$. From the right panel of Fig. 2, we see that $v_{\rm rot}\propto r^{-1/2}$ for $r<100\,\mathrm{pc}$, indicating that $v_{\rm rot}$ decreases with radius. Consequently, the $p$-wave feature, where $\sigma v_{\rm rel.}\propto v_{\rm rel.}^{2}$, results in $\langle\sigma v_{\rm rel.}\rangle\propto r^{-1}$. In addition, $\langle\sigma v_{\rm rel.}\rangle$ increases with $m_{\chi}^{2}$ for $m_{\chi}\gg m_{a}$ as presented in Eq. (19), leading to dash lines ($m_{\chi}=250\,\mathrm{GeV}$) larger than solid lines ($m_{\chi}=25\,\mathrm{GeV}$) by two orders of magnitude. Similarly, those red lines ($m_{a}/m_{\chi}=0.9$) are one order of magnitude smaller than the blue lines ($m_{a}/m_{\chi}=0.1$), because a larger value of $m_{a}/m_{\chi}$ corresponds to a smaller $\langle\sigma v_{\rm rel.}\rangle$.
@@ III.2 Resonance annihilation
where $M_{Z}$ and $M_{W}$ are the masses of the $Z$ and $W$ bosons, respectively. With the assumption of $C_{BB}=C_{WW}$, the partial decay widths of $a\to\chi\overline{\chi}$, $a\to\gamma\gamma$, $a\to ZZ$, and $a\to W^{+}W^{-}$ can be found in detail in Appendix D. The decay mode $a\to\gamma Z$ vanished when $C_{BB}=C_{WW}$.
@@ IV Gamma-Ray Flux from GC
Similarly, in the forbidden DM scenario, BP4 (gold), BP5 (pink), and BP6 (cyan) exhibit spectra in the lower panels. For the forbidden DM scenario with $m_{\chi}=100\,\mathrm{GeV}$, BP4 and BP5 show continuum spectra, whereas BP6 predicts a spectral line, with parameters:
@@ V.2 Constraints of sharp spectral structures from the combined Fermi-LAT and DAMPE data
Thanks to the large statistics, the Fermi-LAT data are widely used in the search of the line-like (Abdo et al., 2010; Ackermann et al., 2015; Bringmann et al., 2012; Liang et al., 2016; Foster et al., 2023; De La Torre Luque et al., 2024; Fan et al., 2024) and box-like (Ibarra et al., 2012; Li et al., 2018; Shen et al., 2021) structures. We use the 15-yr P8R3_V3 CLEAN Fermi-LAT data (Ajello et al., 2021) between 2008 August 4 and 2023 August 3 (Fermi Mission Elapsed Time (MET) from 239557417 to 712715727) to search for sharp spectral structures. To improve the sensitivity of sharp structures, we only adopt the evtype=896 events, which removes $\approx 1/4$ of the data with the worst reconstructed energy quality. The same zenith cut and data-quality filter cut as the previous subsection are applied with Fermitools.
@@ VII Summary and Discussion
Finally, we aim to compare our findings with existing researches in the field. Previous studies Chiang et al. (2020); Alvarez and Yu (2021); Liu et al. (2022); Balaji et al. (2023) focus on $s$-wave annihilation. In contrast, only $a$-resonance annihilation in our work is under $s$-wave condition, highlighting scenario where annihilation cross-sections are enhanced near the BH by reaching the pole velocity. Ref. Cheng et al. (2023b) also examines resonance annihilation through a Higgs portal DM model, while annihilation into photon pairs is not the tree-level contribution. Some studies investigate $p$-wave annihilation near the Sgr A* Shelton et al. (2015); Sandick et al. (2018); Johnson et al. (2019); Cheng et al. (2022) and in the center of M87 Christy et al. (2023). Only Ref. Johnson et al. (2019) examines the axion-portal DM model, but we derive constraints using the halo slope given by S2 data rather than $\gamma\geq 1$. In Appendix E, we show the impact of $\gamma$ on spike halo density and the upper limits of $C_{a\chi\chi}$ in the $a$-resonance scenario. Furthermore, Ref. Johnson et al. (2019) claims that constraints derived from the continuum analyses are more stringent than those from line analyses. However, our analysis demonstrates that integrating both line and continuum analyses using data from Fermi and DAMPE can offer complementary limits. Ref. Shelton et al. (2015) also discusses various search strategies for lines or continuum signals depending on parameter space, but they focus on the Higgs-portal model. In contrast, we demonstrate that the axion-portal model supports a monochromatic line in both $a$-resonance and forbidden DM scenarios.
@@ Acknowledgments
We thank an anonymous referee for useful comments on gravitational potential. YST, MWY, ZQG and ZQX are supported by the National Key Research and Development Program of China (No. 2022YFF0503304), and the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). CTL and XYL are supported by the Special funds for postdoctoral overseas recruitment, Ministry of Education of China (No. 164080H0262403). This research has made use of data obtained from the High Energy Astrophysics Science Archive Research Center (HEASARC), provided by NASA’s Goddard Space Flight Center. This research has made use of the data resources from DArk Matter Particle Explorer (DAMPE) satellite mission supported by Strategic Priority Program on Space Science, and China data service provided by National Space Science Data Center of China.
Appendix A Potential profile
Figure 11: The total gravitational potential $\Phi_{\rm total}$ (illustrated as a red solid line) containing contributions from both BH (black dot-dashed line) and DM (green dot-dashed line, NFW with spike profile). The blue solid line represents the gravitational potential only with the NFW profile contribution.
In Fig. 11, we illustrate the potential functions from BH, DM (NFW with spike profile), NFW profile, and the sum, respectively represented by black dot-dashed line, green dot-dashed line, blue solid line, and red solid line. In the innermost region, the BH gravitational potential dominates, while the NFW DM gravitational potential takes over if $r>10^{-1}\,\mathrm{pc}$.
Appendix B DM self-interaction